### v1

# Measuring the diffuse Galactic synchrotron spectral index and curvature between 45 and 2300 MHz

Abstract
We present an all-sky map of the synchrotron spectral index and curvature between 45 and 2300 MHz at a resolution of 1 calculated from a combination of numerous partial sky empirical measurements. We employ a least-squares parametric fit which relies on removing a free-free emission template and a component separation technique which fits for both synchrotron and free-free emission. We compare our diffuse sky model estimates against those derived from the models widely used in the community (e.g. pysm3 and GSM) employing external datasets that were not included in the estimation process. Our evaluation focuses on identifying the enhanced consistency at both the map level and in pixel-to-pixel correlations, allowing for a more robust verification of our model’s performance. We find our parametric, least-squares synchrotron estimate to be the most reliable across radio frequencies as it consistently provides sky models with average accuracies (when compared to empirical data) of around 20 per cent, whilst other model performances range on average between 10 and 70 per cent accurate. The results obtained have been made publicly accessible online and can be utilized to further develop and refine models of Galactic synchrotron emission.
keywords:
(cosmology:) diffuse radiation – radio continuum: ISM – methods: data analysis
pubyear: 2025pagerange: Measuring the diffuse Galactic synchrotron spectral index and curvature between 45 and 2300 MHzC
1 Introduction
Diffuse Galactic synchrotron emission is caused by charged, relativistic particles, such as cosmic-ray electrons, propagating along the field lines of our Galactic magnetic field (strong). It is the dominant diffuse Galactic emission at frequencies under 1 GHz across the majority of the sky, excluding the central Galactic plane in both intensity and polarization (bennett). Synchrotron emission is of particular interest to several communities due to its predominance in a very wide range of frequencies and because of its complex spectral and spatial behavior. On one hand, for the interstellar medium community measurements of diffuse synchrotron emission places constraints on the Galactic magnetic field strength as well as comic-ray electron propagation (pad21; bracco), provide an understanding of the diffuse background required for transient and supernova remnant detection (trans; erosita) and probe exotic physics such as dark matter annihilation (dm) and outflows/super bubbles from the Galactic center (haze). Within the cosmological community diffuse Galactic synchrotron emission provides a pernicious foreground for both 21 cm intensity mapping and the Cosmic Microwave Background (CMB).
21 cm experiments aim to measure the large angular scale, unresolved, integrated emission from neutral hydrogen atoms across redshift thus mapping the formation of Large Scale Structure in the Universe and determining the redshift of the Epoch of Reionisation. 21 cm experiments are either global experiments which aim to measure the average neutral hydrogen temperature (e.g. EDGES (edges), LEDA (leda), REACH (reach), SARAS (saras)) or intensity mapping experiments which measure the anisotropies around this average (e.g. BINGO (bingo), CHIME (chime), HIRAX (hirax), MeerKLASS (meerklass)). Global and intensity mapping 21 cm experiments measure within the 50-1420 MHz frequency ranges and so suffer contamination from synchrotron emission in intensity and also from polarization through the so-called Polarization-to-Intensity leakage of Stokes parameters Q and U to I (shaw14). Similarly, the constraints of cosmological parameters as well as the detection of the faint $B$-mode polarization signal from CMB data is strongly contaminated by Galactic foregrounds, if not properly taken into account and/or removed from the CMB intensity and polarization maps (so; litebird; planckPS).
The synchrotron emission spectrum at each pixel on the sky ($p$) is modeled as a power law,
$T_{\nu}(\nu,p)=T_{\nu_{0}}(\nu,p)\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{\text{s}}(p)},$ (1)
where the synchrotron spectral index $\beta_{\rm s}$ is measured between the frequencies $\nu$ and $\nu_{0}$. $T_{\nu_{0}}$ is a synchrotron temperature map at a particular frequency. As synchrotron emission decreases with increasing frequency, the lowest frequency, highest resolution all-sky radio map available is often considered to be the best proxy for a synchrotron emission amplitude template. Since its inception, in the early 1980s, the Haslam 408 MHz (haslam) all-sky map at a resolution of 56 arcmin has universally been used to represent a full sky view of synchrotron emission. It is well understood that the synchrotron spectral index should change within different lines of sight in the sky, due to energy losses as the charged relativistic particles travel through the Galaxy. A widely accepted measurement of the synchrotron spectral index across the full sky was made in mamd, where the authors explored several methods to very different effect. All the methods involve trying to form an all-sky map at 5 resolution between 408 MHz and 23 GHz using Haslam and WMAP (wmap) data. The mamd spectral index has been vastly employed, particularly in modelling synchrotron emission in the frequency regime of typical CMB experiments (20-100 GHz). In fact, the latest models of the Python Sky Model pysm3 suite, pysm3, employ the mamd template together with the recently available S-PASS data (Krachmalnicoff_2018). Moreover, the limit in resolution has been overcome by artificially adding small scales in the spectral index map (pysm3).
Not only is the spectral index expected to change spatially, it is also believed to vary across frequency. Throughout the early 2000s an increasing number of researchers were noticing a discrepancy between the spectral index measured at radio frequencies by ground-based telescopes and microwave measurements made by satellites. deOliveira2008 note that the spectrum of synchrotron emission seems to steepen with increasing frequency after a consideration of over twenty partial sky maps between 10 MHz and 100 GHz. The ARCADE2 balloon-borne experiment (arcade) measured a region of the sky between 3 and 90 GHz and fit an empirical form for the believed change in spectral index over frequency, denoted as spectral index ‘curvature ($c_{s}$)’:
$\beta(p)=\beta_{0}(p)+c_{s}\,{\rm{ln}}\left(\frac{\nu}{\nu_{0}}\right),$ (2)
[…]
Simulations of diffuse Galactic emissions have particular use within the 21 cm and CMB communities as they provide a realistic test bed for cosmologists to test their component separation methods which enable the recovery of a cosmological signal of interest from an empirical data set which contains Galactic foregrounds. Additionally, instrumental calibration schemes may require an initial estimate of the radio sky temperature and so an understanding of synchrotron emission both spatially and spectrally is required. Total diffuse emission models, such as the Global Sky Model (GSM (gsm)), pysm3 and Bayesian Global Sky Model (bayes), are especially useful as they provide temperature estimates for the combined diffuse Galactic emission temperature (which will be some combination of synchrotron, free-free, anomalous microwave and thermal dust emission) per pixel across a large frequency for the whole sky.
Lately, the community has been focusing on improving the synchrotron models (cnnBeta; synax; tigress). This involves accurately accounting for the emission spatial and spectral variations (e.g. by incorporating the curvature parameter), and by overcoming the limitation of low angular resolution synchrotron templates. The objective of this study is to tackle both these two goals by combining information from a total of 36 distinct partial and all-sky maps, each of which are available at varying resolutions. This integration is aimed at generating comprehensive maps that represent an updated template of the diffuse Galactic synchrotron spectral index and curvature intended to be employed across a wide range of frequencies, specifically from 45 to 2300 MHz.
The paper is organized as follows: in section 2 we introduce the publicly available data and section 3 details the processing steps required to combine the data sets into a parametric fit of the spectral index and curvature. We explore two ways to perform the fit, a least-squares parametric fit which relies on removing a free-free emission template and a component separation technique which fits for both synchrotron and free-free emission. In section 4 we form an all-sky model of synchrotron emission by scaling the Haslam data using the fitted spectral index and curvature maps. Lastly, in section 5 we present our conclusions.
2 The data
In this work we make use of publicly available radio surveys that cover both the northern and southern hemispheres. Table 1 summarizes the main observational details of the publicly available surveys employed within this work: Maipu/MU (mu), LWA1 (lwa), OVRO-LWA (ovro), Landecker (land), Haslam (haslam), GMIMS-HBN (hbn) and GMIMS-STAPS (staps). For the 408 MHz data we make use of the destriped map reprocessed by rem.
Survey Freq. (MHz) Res. () $\sigma_{{\rm{cal}}}$ ($\%$) $\sigma_{{\rm{t}}}$ Coverage
Maipu/MU 45 5.0 7.0 0.23b Full
OVRO-LWA 41.8 0.29 5.0 0.8a North
OVRO-LWA 47.0 0.27 5.0 0.8a North
OVRO-LWA 52.2 0.25 5.0 0.8a North
OVRO-LWA 57.5 0.25 5.0 0.8a North
OVRO-LWA 62.7 0.25 5.0 0.8a North
OVRO-LWA 67.9 0.25 5.0 0.8a North
OVRO-LWA 73.2 0.25 5.0 0.8a North
LWA1 50 3.3 5.0 16a North
LWA1 60 2.7 5.0 10a North
LWA1 70 2.3 5.0 7a North
LWA1 74 2.2 5.0 6a North
LWA1 80 2.0 5.0 5a North
Landecker 150 5.0 5.0 0.62b Full
Haslam 408 0.93 10.0 0.1b Full
GMIMS-HBN 1383 0.67 8.0 0.02b North
GMIMS-HBN 1418 0.67 8.0 0.02b North
GMIMS-HBN 1456 0.67 8.0 0.02b North
GMIMS-HBN 1487 0.67 8.0 0.02b North
GMIMS-HBN 1499 0.67 8.0 0.02b North
GMIMS-HBN 1521 0.67 8.0 0.02b North
GMIMS-HBN 1614 0.67 8.0 0.02b North
GMIMS-HBN 1625 0.67 8.0 0.02b North
GMIMS-HBN 1660 0.67 8.0 0.02b North
GMIMS-HBN 1700 0.67 8.0 0.02b North
GMIMS-HBN 1712 0.67 8.0 0.02b North
GMIMS-STAPS 1324 0.33 10.0 0.016b South
GMIMS-STAPS 1349 0.33 10.0 0.016b South
GMIMS-STAPS 1374 0.33 10.0 0.015b South
GMIMS-STAPS 1456 0.33 10.0 0.013b South
GMIMS-STAPS 1524 0.33 10.0 0.012b South
GMIMS-STAPS 1609 0.33 10.0 0.011b South
GMIMS-STAPS 1628 0.33 10.0 0.010b South
GMIMS-STAPS 1700 0.33 10.0 0.009b South
GMIMS-STAPS 1749 0.33 10.0 0.008b South
GMIMS-STAPS 1770 0.33 10.0 0.008b South
Table 1: The datasets employed in this work; survey name, central frequency, angular resolution, calibration errors, instrumental noise level and sky coverage. Instrumental noise ($\sigma_{{\rm{t}}}$) is given either in Jy per beam (a) or K ($b$).
3 Methods
Total emission ($T_{\rm{tot}}$) measurements of the sky (per pixel $p$) across radio frequencies ($\nu$) contain several temperature contributions:
$T_{\rm{tot}}(p,\nu)$ $=T_{\rm{diff}}(p,\nu)+T_{\rm{ps}}(p,\nu)+T_{\rm{cosmo}}(p,\nu)+T_{\rm{cosmo}}(\nu)$ (3)
$+T_{\rm{inst}}(\nu),$
where $T_{\rm{diff}}$ is the contribution from diffuse Galactic emissions, $T_{\rm{ps}}$ is from point sources large enough to be measured as distinct features by the instrumental beam and $T_{\rm{cosmo}}$ is from cosmological backgrounds such as the CMB, the neutral hydrogen background (21 cm emission) and the unresolved radio point source collectively known as the Cosmic Radio Background (CRB). Cosmological backgrounds contribute a constant single value across all pixels of the sky, such as the CMB monopole of 2.725 K as well as anisotropies around this monopole value, which vary across pixels. $T_{\rm{inst}}$ represents a frequency dependent monopole inherent to each individual experiment. Whilst, all of the observational data used in this work are calibrated from receiver units into Kelvin, only one survey has an absolutely calibrated temperature zero-level. This means that all other surveys are measuring temperature variations across the sky above an arbitrary zero-level determined by their own data reduction pipeline.
In order to measure the diffuse Galactic synchrotron emission spectral index, we must isolate the diffuse Galactic synchrotron emission at each pixel and frequency. For this we make two main assumptions about radio data:
  1. 1.
    Diffuse Galactic emission at frequencies under 3 GHz consists of only synchrotron and free-free contributions. Anomalous microwave and thermal dust emissions are negligible within this frequency range.
  2. 2.
    The CMB, CRB and 21 cm spatially varying signals at each frequency are so small in comparison to the spatially varying signal coming from our own Galaxy at each frequency that they can be considered negligible.
Rewriting Eq. 3 with these assumptions in mind yields:
$T_{\rm{tot}}(p,\nu)=T_{\rm{sync}}(p,\nu)+T_{\rm{ff}}(p,\nu)+T_{\rm{ps}}(p,\nu)+T_{\rm{back}}(\nu),$ (4)
where $T_{\rm{sync}}(p,\nu)$ and $T_{\rm{ff}}(p,\nu)$ represent diffuse Galactic synchrotron and free-free emission, respectively and $T_{\rm{back}}(\nu)$ is the combination of any monopole temperatures present. All that is now required to fit a synchrotron spectral index across the different surveys is to 1) calculate and subtract the frequency-dependent zero levels, 2) model and remove the diffuse free-free emission and 3) mask out and inpaint the point source contributions.
3.1 Zero-Level calibration
In general fewer surveys with absolute zero-level calibration exist than those without, as an additional reference black-body emitter is required within the instrumentation to provide regular zero-level checks. However, it is possible to measure a survey zero-level through comparison with another, zero-level calibrated survey (jonas).
Both free-free and synchrotron emission can be modeled as a power law. If an observation within a single pixel across two frequencies represents pure synchrotron then:
$T_{\rm{sync}}(p,\nu_{2})=T_{\rm{sync}}(p,\nu_{1})\times\left(\frac{\nu_{2}}{\nu_{1}}\right)^{\beta_{\rm sync}}$ (5)
If the emission at both frequencies is free-free emission than the spectral index calculated is the free-free spectral index and if the emission present is some combination of both free-free and synchrotron emission than the spectral index represents the scaling across frequency of that mixture.
The total diffuse emission measured will include the survey zero-levels:
$T_{\rm{diff}}(p,\nu_{2})+c_{v2}=\left(T_{\rm{diff}}(p,\nu_{1})-c_{v1}\right)\times\left(\frac{\nu_{2}}{\nu_{1}}\right)^{\beta_{\rm diff}}$ (6)
The Haslam 408 MHz all-sky survey has both a temperature scale and zero-level calibration and, thanks to its widespread use, the monopole due to the combined CMB, CRB and 21 cm emission has been calculated as $8.9\pm 1.3$ K (wehus). Therefore for any survey compared with the Haslam survey:
$T_{\rm{diff}}(p,\nu_{2})=(T_{\rm{tot}}-8.9)(p,408)\times\left(\frac{\nu_{2}}{408}\right)^{\beta_{\rm diff}}-c_{v2}$ (7)
As long as the maps at both frequencies are observing the same emission then the above equation can be fit using linear regression providing a value for $c_{v2}$ as the fitted y-intercept. If, however, the difference in frequency means that one survey only observes synchrotron emission whilst the seconds survey observes synchrotron and non-negligible free-free emission then the two datasets will have a poor correlation coefficient and the linear regression will fail.
The zero-level value for each frequency map was calculated by:
  1. 1.
    Smoothing all maps to a common resolution assuming Gaussian beams.
  2. 2.
    Each map at a given frequency $\nu$ was subdivided into sub-regions corresponding to HEALPix pixels at $N_{{\rm{side}}}=8$ (gorski). Each sub-region encodes 1024 pixels.
  3. 3.
    Linear regression was performed within each sub-region between the considered frequency $\nu$ and 408 MHz after having removed the 8.9 K offset from the Haslam map.
  4. 4.
    The mean of all the region offsets was used as the $\nu$ zero-level, with the standard error on the mean providing the offset error.
The errors associated with each dataset are just the survey calibration errors except for the Haslam data where the total error ($\sigma_{\rm{haslam}}$) is calculated as:
$\sigma_{\rm{haslam}}=\sqrt{{\sigma_{\rm{cal}}}^{2}+{\sigma_{\rm{offset}}}^{2}},$ (8)
where $\sigma_{\rm{offset}}$} is 1.3 K.
Figure 14 displays the offset values calculated within each sub-region for the single-dish empirical data at each frequency. Figure 13 displays these offsets as a histogram distribution. The mean offset was used as the final map offset, with the standard error on the mean providing the uncertainty. Note that we do not calculate the zero-level offset for the OVRO-LWA data as they are interferometric data and therefore missing any angular scales larger than the shortest array baselines can measure. The calculated offsets for each single-dish observation are listed in Table 3.
3.2 Removal of free-free template and least-squares fitting
After the offset removal, each map contains some mixture of both diffuse free-free and diffuse synchrotron emission, as well as the point source emission contribution. In order to isolate pure diffuse synchrotron emission we must either remove a free-free emission template or specifically fit for the synchrotron emission. We explored both these options; this section describes the first approach which removes a free-free template and then performed a least-squares fit to the remaining emission (assumed to be pure synchrotron).
Figure 1: Top: Emission measure template used in this work. Bottom: Electron temperature template used in this work.
Free-free emission can be parametrized using electron temperature ($T_{e}$) and emission measure (EM) as follows:
$T_{{\rm{ff}}}=T_{e}\,(1-e^{-\tau}),$ (9)
where
$\tau=0.05468\,T_{\mathrm{e}}^{-3/2}\,\nu_{9}^{-2}\,{\rm{EM}}\,g_{\rm ff},$ (10)
where
$g_{\rm ff}(\nu,T_{\mathrm{e}})=\ln\left\{\exp\left[5.960-\frac{\sqrt{3}}{\pi}\ln(\nu_{\rm GHz}T_{4}^{-3/2})\right]+e\right\},$ (11)
and $\nu_{9}$ and $T_{4}$ are $\nu={10^{-9}}$ Hz and $T_{\mathrm{e}}=10^{-4}$ K respectively.
Figure 1 presents the emission measure template and electron temperature template used in this work to simulate free-free emission at each total emission map frequency. We employ the huts EM map, obtained with a joint inference model of the Planck (ffparam) and H-$\alpha$ (fink) emission measure maps. huts publicly released the EM map at a 6 arcmin resolution and Nside 256 . The $T_{e}$ map we employ was produced through the application of the Bayesian component separation framework Commander of the Planck PR2 (ffparam), available at a resolution of $1^{\circ}$ and Nside 256. Figure 2 illustrates the typical contribution of a free-free emission (from our model) to the diffuse Galactic emission by plotting the temperature RMS across the single dish empirical observations smoothed to 5 resolution.
Figure 2: RMS of maps as a function of frequencies estimated in the common footprint observed at 5 degree resolution.
Figure 3: An example of the per-pixel fit (Galactic latitude and longitude of the pixel noted on the figure) performed to determine the synchrotron spectral index and magnitude of spectral index curvature. Empirical data points are given in blue; the best fit to the data is the orange line.
After the offset and free-free emission removal the empirical maps contain only diffuse Galactic synchrotron and point source emissions. We choose to leave the point sources in the maps for the joint fit and to only remove them in the final fitted parameter maps. The logic behind this was driven by the fact that one map viewing the sky at a resolution of 56 arcmin will see the point sources smoothed with a 56 arcmin FWHM beam thus spreading some of the point source temperature into the diffuse emission temperature. When we smooth a second, finer resolution map to 56 arcmin to compare with the first map we also want the point source temperature contribution to be smoothed into the diffuse temperature contribution as that is the most faithful representation of the sky as measured through a 56 arcmin beam. Therefore we choose to ignore the points sources for the fit in favor of masking them out in the final parameter maps.
Assuming pure diffuse Galactic synchrotron emission the fitted model becomes:
$\frac{T_{\rm{sync}}(p,\nu)}{T_{\rm{sync}}(p,\nu_{0})}=\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{\rm sync}},$ (12)
where
$\beta_{\rm sync}=\beta_{\nu_{0}}+c\ln(\frac{\nu}{\nu_{0}}),$ (13)
and the two free parameters are the spectral index ($\beta_{\rm sync}$) and curvature ($c$). We choose $\nu_{0}$ to be the lowest frequency available from the empirical data as the synchrotron emission in this map has the highest signal to noise ratio due to the negative spectral index of synchrotron emission.
A per-pixel fit of empirical data can only be performed when each pixel represents the same angular information of the sky so each empirical map must be smoothed to a common resolution and downgraded to the same Nside. We perform four fits across four different data combinations. This is necessary as several of the empirical maps are northern or southern hemisphere only, some data sets are single-dish and have a resolution limit and other data sets are interferometric and are missing large scale information on the sky. The four combinations are:
  1. 1.
    Northern coarse (NC): LWA1 45, 50, 60, 70, 74, 80  MHz, Haslam 408 MHz and GMIMS 1383, 1418, 1456, 1486, 1499, 1521, 1614, 1625, 1660, 1700, 1712 MHz data. All maps were smoothed to a common resolution of 5 and downgraded to Nside 256. All instruments in this set are single-dish and so no angular scales over 5 are missing. Figure 9 shows the empirical data employed for the coarse northern fit.
  2. 2.
    Southern coarse (SC): Maipu/MU 45 MHz, Haslam 408 MHz, STAPS 1324, 1349, 1374, 1456, 1524, 1609, 1628, 1700, 1749, 1770 MHz data. All maps were smoothed to a common resolution of 5 and downgraded to Nside 256. All instruments in this set are single-dish and so no angular scales over 5 are missing. Figure 10 shows the empirical data employed for the coarse southern fit.
  3. 3.
    Northern fine (NF): OVRO-LWA 41, 46, 52, 57, 62, 67, 73 MHz, Haslam 408 MHz and GMIMS 1383, 1418, 1456, 1486, 1499, 1521, 1614, 1625, 1660, 1700, 1712 MHz data. All maps were smoothed to a common resolution of 1 and downgraded to Nside 256. Figure 11 shows the OVRO-LWA data employed for the fine northern fit. This fit is only used to provide the 1 to 5 angular resolution information e.g the finer scaler anisotropies around the large scale patterns identified within the northern coarse fit.
  4. 4.
    Southern fine (SF): Haslam 408 MHz and STAPS 1324, 1349, 1374, 1456, 1524, 1609, 1628, 1700, 1749, 1770 MHz data. All maps were smoothed to a common resolution of 1 and downgraded to Nside 256. The lack of high resolution public data covering the southern hemisphere at frequencies lower than 408 MHz results in a poorly constrained fit and so instead leaving the spectral index and amount of curvature to be completely free parameters, they were constrained at each pixel to be within $\pm 5$ per cent of the value obtained in the Southern coarse fit. Figure 12 shows the empirical data employed for the fine southern fit. This fit is only used to provide the 1 to 5 angular resolution information e.g the finer scaler anisotropies around the large scale patterns identified within the southern coarse fit.
We find the least-squares per-pixel fit performs best (with the lowest reduced $\chi^{2}$) within log-space and with an additional free parameter ($f$) providing a multiplicative factor:
$\frac{T_{\rm{sync}}(p,\nu)}{T_{\rm{sync}}(p,\nu_{0})}=f\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{\rm sync}},$ (14)
which, providing our model is adequate, should be very close to 1, or very close to 0 in log-space. This parameter is ancillary to help the fitting algorithm in dealing with pixels that contained systematic errors not being taken into account within our model, which is purely motivated by synchrotron emission behavior. Figure 3 shows an example of a per-pixel fit for the Northern coarse data set; the fitted parameters and reduced $\chi^{2}$ are given on the plot.
3.3 Component separation
In this section we present the complementary approach to estimate the synchrotron spectral parameters estimated through a parametric component separation with fgbuster (fgsoft; Stompor_2009; Stompor_2016).
fgbuster is commonly adopted in recovering CMB data from the multi-frequency observations $d_{\nu}$ of multiple astrophysical components $s$, i.e. the CMB together with the Galactic foregrounds like thermal dust and synchrotron. Each signal scales with frequency with the so-called mixing matrix, $A(\beta)$ depending on a set of spectral parameters, $\beta$:
$d_{\nu}=A({\beta})s+n_{\nu}$
with $n_{\nu}$ being the instrumental noise of each frequency map. The methodology essentially consists in finding the optimal set of parameters $\beta$ and of component maps $s$, that maximizes the spectral likelihood $\mathcal{L}$:
$2\ln\mathcal{L}(s,\beta)=(d-A(\beta)s)^{T}N^{-1}(d-A(\beta)s)+\mathrm{const.},$
with $N^{-1}$ being the inverse noise covariance matrix.
As the interstellar medium properties changes, the value of the Galactic spectral parameters are expected to spatially vary along multiple lines of sight. This further complicates the recovery of the CMB polarized signals, as it might leave a mis-modeling bias into the faint B-mode signal if those spatial variations are not properly taken into account. However, latest results in the literature (Errard_2019; Puglisi_2022; rizzieri2025), have shown that the quality of the recovered signals significantly improves both in terms of convergence to the solution and in the reduced level of mis-modelling bias when the spectral likelihood is estimated on a subset of pixels. Practically, these sub-group of pixels are optimally chosen to be large enough to increase the SNR of each subset and small enough to reduce the mis-modelling bias.
In this work, we employ fgbuster for the first time to fit for both synchrotron and free-free emissions, the former modelled as power-law with a non-zero running of the spectral index:
$s_{\rm{sync}}(p,\nu)\propto\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{s}(p)+c_{s}(p)\ln\ {\nu}/{\nu_{c}(p)}},$ (15)
where we explicitly express the dependence of the emission by a given pixel $p$, $\nu_{0}=408$ MHz is the reference frequency, whereas $\nu_{c}$ is a parameter to be fitted and encodes the pivotal frequency for the $c_{s}$ parameter. generally set to be $\nu_{c}\neq\nu_{0}$. Free-free emission is modelled as :
$s_{\rm{free}}(p,\nu)\propto\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{f}(p)},$ (16)
where $\beta_{f}$ is the spectral index, and $\nu_{0}=408$ MHz also in this case.
Therefore, the set of parameters fit by fgbuster are thus : $\beta=\left[\beta_{s},c_{s},\nu_{c},\beta_{f}\right]$ and are assumed to be spatially variable in the sky.
[…]
We run separately fgbuster for the 4 datasets described in subsection 3.2, i.e. encoding the North and South coarse and fine resolution maps. Each run was performed on a local machine and took 50 minutes for the North maps, and 30 min for the South ones, (as the former dataset presents a larger footprint than the latter).
3.4 Combination and point source removal
The four different parameter fits result in four different maps for both the synchrotron spectral index and the spectral index curvature at the coarse and fine angular resolutions. It should be noted that as the value for $\beta_{s}$ changes across frequency in the model used in this work, we present maps of $\beta_{s}$ that are relative to the chosen values of $\nu_{0}$: 45 MHz for the least-squares fit and 408 MHz for fgbuster.
The large-scale fits were performed at $5^{\circ}$; the OVRO-LWA data were not included in this fit as they lack angular information on the sky larger than several degrees. The finer angular information in the northern hemisphere was provided by the OVRO-LWA data. The community are currently lacking the equivalent, high resolution, low frequency data covering the southern hemisphere. As the northern fit contains interferometric data, missing large angular scales and the southern fit only covers the frequency range 408 to 1170 MHz the 1 parametric fit is only used to constrain the spectral index and curvature anisotropies. The coarse northern and southern fit were used to provide all the angular information above 5. We then joined together each north and south parameter maps. As we had more frequency data available in the Northern surveys, we prioritized the northern results over the southern ones when considering the overlap region between the two. Regions which were not observed at all within the northern fit were supplemented with corresponding data from the southern maps.
Figure 4: Combined maps of NC,NF, SC and SF synchrotron spectral parameters (left panel) $\beta_{s}$ and (right panel) $c_{s}$ obtained with (top row) a parametric fit and (bottom row) with fgbuster . Note that $\beta_{s}$ is calculated at 45 MHz for the parametric fit and at 408 MHz for fgbuster .
Figure 5: Angular power spectra estimated from (left panel) $\beta_{s}$ and (right panel) $c_{s}$ full-sky maps shown in Figure 4. The spectra obtained from the parametric fit and fgbuster maps are respectively shown in (solid orange) and (solid blue), in (solid black) we also show the power spectrum estimated from the template of the pysm3 s7 model.
The joined maps lacked temperature information around the North and the South Celestial poles. Additionally, we wished to exclude extragalactic points source from our analysis. For the point source removal we made use of the Second Planck catalog of compact sources (planckPS) at 30 GHz. Only sources with galactic latitudes ($b$) further than $7^{\circ}$ from $b=0^{\circ}$ were excluded from the parameter maps to ensure that only extragalactic sources were removed. We filled these missing/masked locations with the diffusive inpainting algorithm presented in Puglisi_2020. The algorithm operates by filling in the pixels with the mean value, which is calculated based on its closest neighboring pixels. This process is iteratively carried out, until the 2-norm difference of maps from two consecutive iterations reaches an absolute tolerance of $10^{-4}$. For the maps shown in this work, less than 700 iterations are needed to achieve the set tolerance.
Once both NC-SC and NF-SF maps were clipped together and inpainted we combined them into one single map employing the spherical harmonic decomposition routines available in the HEALPix library. We then low-(high-)pass filtered the coefficients from the coarse $a_{\ell m}^{C}$ (fine $a_{\ell m}^{F}$ ) maps as follows :
$a_{\ell m}^{tot}=a_{\ell m}^{C}\sigma_{\ell}+a_{\ell m}^{F}\left(1-\sigma_{\ell}\right)~~~,$ (17)
with $\sigma_{\ell}$ being a sigmoid filter:
$\sigma_{\ell}=\left[1+e^{(\ell-\ell_{0})/w)}\right]^{-1}~~~,$ (18)
with $\ell_{0}=40$ and $w=2$ governing respectively the scale and the width of the filter as a function of multipoles.
The top row of Figure 4 shows the combined northern and southern, $1^{\circ}$ resolution synchrotron spectral index and curvature maps with point sources and missing data inpainted for the least-squares parametric fit. The spectral index at 45 MHz generally fits the expected range (3.2-3.2 to 2.2-2.2) for synchrotron emission, with the exception of those few pixels which present indices flatter than 2.1-2.1. A spectral index between 2.1-2.1 to 2.0-2.0 is generally expected for free-free emission and any values flatter than that are likely to just be associated with a poor fit. Whilst the reduced $\chi^{2}$ maps for each of the four individual fits (northern coarse and fine, southern coarse and fine displayed in Figure 15 and Figure 16,) shows a good fit over all the pixels it is clear from the spectral index and curvature maps that these two parameters are highly degenerate and display a strong anti-correlation. The red areas within the spectral index map display a unusually flat spectral index and these exact same areas are shown in blue in the curvature map as they display an usually large degree of curvature. The full covariance matrices of correlated errors are more revealing when determining the accuracy of the fit than just the reduced $\chi^{2}$ maps alone. These matrices are calculated as part of the fitting and so will be available within the software release associated with this work.
Figure 6: Fractional difference maps between the empirical data and the models at 159, 820, 2303, 2326 and 11000 MHz from left to right respectively. The top row shows the differences for the parametric model, the middle for the GSM and the bottom row for the pysm3.
In Figure 17 we show the spectral parameters obtained with fgbuster runs for the combined maps at both the low and high angular resolution. The final inpainted maps are shown in the bottom panel of Figure 4. We observe that the $\beta_{s}$ estimates are systematically steeper than those obtained from the parametric fit, particularly in regions near the the Galactic plane and along the spur. A similar trend is evident in the $c_{s}$ map, where one can clearly distinguish the regions in which the northern and southern surveys are included. As indicated by strong correlation of $\beta_{s}$ and $c_{s}$ maps and by the large values (>10>10) of the reduced $\chi^{2}$, we can argue that the uncertainties have likely been underestimated in the fgbuster runs, This, in turn, introduces systematic errors that distort the inferred values of the spectral parameters, leading to biases and unreliable estimates.
Finally, we estimate the power spectrum of angular correlations from $\beta_{s}$ and $c_{s}$ maps. We remark here that our maps employ large and small scales together. Moreover, the diffusive inpainting involved few percentage of the sky allowing us to perform a proper Spherical Harmonic decomposition and to mitigate the impact of extra-galactic sources. The angular power spectra [^1] Estimated with Healpix routines https://healpy.readthedocs.io/en/latest/healpy_spht.html obtained from the parametric least-squares fit and fgbuster maps are shown in Figure 5 and are compared with the ones estimated from the template of the pysm3s7’ model, which uses the Haslam map as a synchrotron amplitude template, the model 4 synchrotron spectral index map from mamd rescaled using S-PASS data and a spatially constant value for the spectral index curvature taken from kogut. We notice that the three $\beta_{s}$ power spectra follow almost a similar power law scaling, although the spectrum estimated from the pysm model is an order of magnitude lower at lower multipoles ($\ell\lesssim 10$). The power spectra obtained with fgbuster and the parametric fit instead clearly present the effects of having combined two maps with different angular resolutions, e.g. the loss of power at around $\ell\sim 40$ and $\ell>200$, corresponding respectively to the beam angular size of 5 and 1 $\deg$. Being the fgbuster data products pixellated at $N_{\rm side}=64$,the power spectra of Figure 5 in solid blue have been evaluated at $\ell_{max}=3N_{\rm side}-1$. The $c_{s}$ power spectra obtained with fgbuster and with the parametric fit show comparable amplitude, about 2 orders of magnitude larger than the s7 model. Although the two maps have almost similar full-sky average values:0.048,0.073,0.043-0.048,-0.073,-0.043, respectively for the parametric, fgbuster and s7 model, the discrepancy is due to a larger variance of the estimated $c_{s}$ maps with respect to the pysm3 one. We remark here that this is somewhat expected as the s7 has been derived from the Haslam synchrotron template matching the measured values reported in kogut and has not been estimated from a fitting procedure.
4 Results
The synchrotron spectral index and curvature parameter can be used alongside the free-free template employed in this work to construct an all-sky diffuse emission model below 2 GHz as follows:
$T_{\rm{diffuse}}(p,\nu)=T_{\rm{sync}}(p,\nu)+T_{\rm{ff}}(p,\nu),$ (19)
where $T_{\rm{ff}}(p,\nu)$ is calculated as in subsection 3.2 and
$T_{\rm{sync}}(p,\nu)=A_{408}(p)\left(\frac{\nu}{\nu_{0}}\right)^{\beta_{s}(p)+c_{s}(p)\ln({\nu}/{45})},$ (20)
where $A_{408}$ is the synchrotron amplitude template at 408 MHz provided by the 56 arcmin Haslam data and $\beta_{s}$ and $c_{s}$ are the fitted spectral index and curvature map at $1^{\circ}$ resolution.
159 MHz 820 MHz 2303 MHz 2326  MHz 11000 MHz
Parametric 14 - 39 9 - 29 13 - 39 9 - 32 21 - 100
GSM 12 - 57 172 - 343 17 - 39 6 - 24 18 - 47
pysm3 22 - 58 18 - 50 34 - 67 36 - 60 57 - 74
fgbuster 30 - 90 23 - 48 57 - 80 60 - 77 91 - 103
Table 2: Percentage difference ($25^{\rm{th}}$ and $75^{\rm{th}}$ percentile) between sky models and empirical data at five frequencies. The empirical maps cover different regions of the sky but the Galactic plane is always excluded from the difference calculation.
Figure 7: First to last row: correlation and histogram comparison between various sky models and empirical partial sky data at 159, 820, 2303, 2326, 11000 MHz, respectively. The scatter plots exclude the Galactic plane region.
To validate our sky model we enlist the use of public partial sky data which were not used within our fit: EDA2, Dwingeloo, Rhodes, S-PASS and QUIJOTE. The EDA2 survey is a southern survey at 159 MHz with a resolution of $3.1^{\circ}$ (EDA2), Dwingeloo surveyed the northern hemisphere at 820 MHz with a resolution of $1.2^{\circ}$ (dwing), the Rhodes southern hemisphere survey (jonas) observed at a resolution of 20 arcmin at 2326 MHz, S-PASS conducted a 2303 MHz survey with a resolution of 8.9 arcmin (spass) and QUIJOTE MFI are observing the northern sky at 11, 13, 16 and 18 GHz with a resolution of $1.0^{\circ}$ (q11).
We compare our diffuse sky model estimates with those of the GSM and pysm3. For the GSM we use the high resolution model which produces sky estimates at 48 arcmin; as the GSM represents total diffuse emission we do not need to include anything else. This is the same for the fgbuster sky model estimates which already include both synchrotron and free-free emission. For the pysm3 we use the ‘s7’ synchrotron model. Both the pysm3 and the least-squares parametric fit give synchrotron emission models so we add to them the free-free emission template (scaled to the appropriate frequency) from subsection 3.2. The model/data comparisons are performed at resolutions of $3.1^{\circ}$ for the EDA2 data, $1.2^{\circ}$ for the Dwingeloo data and at $1.0^{\circ}$ for the Jonas, S-PASS and QUIJOTE data.
Figure 8: $c_{s}$ (left) and $\sigma(c_{s})$ maps obtained with the parametric least-squared fit shown within the ARCADE2 observational footprint (kogut) and in Orthografic projection centered in $\ell,b=(60^{\circ},0^{\circ})$. For the sake of comparison, we set the color-scale of the $c_{s}$ map to be the same as the one in (kogut, Fig.4).
Figure 6 shows the fractional difference maps between the diffuse sky models and the data defined as:
$\delta=\frac{m_{\rm data}-m_{\rm model}}{m_{\rm data}},$
with $m_{\rm data}$ and $m_{\rm model}$ being respectively the map observed and the one obtained by evaluating Equation 20 from the fit spectral parameters at the same frequency as the data. In order for the empirical data to only contain diffuse and compact emissions the constant map zero-level was calculated and removed using the technique described in subsection 3.1. The difference maps are arranged by frequency from left to right (159, 820, 2303, 2326 and 11000 MHz) and by model from top to bottom (Parametric, GSM, pysm3 and fgbuster). Interestingly each model struggles to replicate the empirical data across different regions of the sky as frequency changes; there is no consistent problem area, e.g. as in the Galactic plane. The GSM provides a good sky model except at 820 MHz. In particular, we anticipate that the s7 model implemented in pysm3 would have exhibited larger residuals when compared against data at frequencies $1\,\mathrm{GHz}$. This is because pysm3 specifically validated the reliability of this model only for frequencies above this threshold, implying that its performance may degrade when extrapolated to lower-frequency regimes. This could explain the reason of larger residuals at 159 GHz. The parametric model performs well at all frequencies except at 11000 MHz at high galactic latitudes; this motivates our decision to recommend the parametric sky model for estimating the sky between 45 and 2300 MHz. The fgbuster model can be seen to overestimate the sky at low frequencies and underestimate it at higher frequencies.
In Figure 7 we further quantify the differences between the empirical data and sky models through the use of correlation plots and histogram distributions. As sky models are most typically used outside of the Galactic plane within the fields of CMB and 21 cm cosmology we choose to mask out the Galactic plane from some of our results. This is done through the use of the Planck GAL080 mask  [^2] https://pla.esac.esa.int which masks out the Galactic plane, leaving 80 per cent of the sky pixels unmasked. The left-hand column shows correlation plots between the data and models outside of the Galactic plane mask. The correlation plots feature a black line of gradient 1 to visually highlight the slope a perfect model would have. The parametric model has the highest correlation coefficient with the data at 159 and 820 MHz whilst the GSM has the highest correlation coefficient at 2303, 2326 and 11000 MHz. The middle and right-hand columns show histogram distributions of the absolute percentage differences between the model and data for all four models, in the middle column the full available sky is used whereas the Galactic plane is masked out for the right-hand column plots. The median absolute percentage differences are marked on the histograms using dotted lines. Table 2 is included to aid the interpretation of the histograms in Figure 7 as it shows the 25th and 75th percentiles between the data and the four models when the Galactic plane has been excluded from the calculation. At 159, 820 and 2303 MHz the parametric model provides lowest median difference between the data and the empirical model whilst at 2326 and 11000 MHz the GSM provides the lowest median difference between the data and the empirical mode.
We finally compare the curvature parameter with the one estimated by kogut. They combined radio surveys at 22 MHz, 1.4, 3 and 10 GHz, and observed a steepening in the radio spectrum corresponding to a curvature parameter of $c_{s}=-0.052\pm 0.005$ as estimated within the ARCADE2 footprint. Figure 8 shows the estimated $c_{s}$ obtained from the parametric fit. Therefore, within the observational patch defined by kogut, we compute the mean value of $c_{s}$ using weights given by the inverse of the total uncertainty, $\sigma(c_{s})$, where $\sigma(c_{s})$ is obtained as the quadratic sum of the errors from the coarse and fine estimates (see Figure 15 and Figure 16). We obtain $c_{s}=-0.0517\pm 0.0007$, compatible, albeit more accurate, with the value reported by kogut.
5 Conclusions
[…]
These results were obtained through a parametric fit of 36 empirical sky maps. These maps differed in sky coverage, namely northern and southern hemisphere, as well as angular resolution. Only two of the maps used were made using instrumentation capable of measuring the absolute zero-level. This work was made possible thanks to the wealth of high quality radio data that has been made public, enormous progress in the production of free-free emission templates and the analytical power of a simple linear regression technique: the temperature-temperature plot. Four separate fits were performed: the northern 5, the southern 5, the northern 1 and the southern 1 fits. The northern and southern hemispheres were joined together, favoring the northern data in the overlap region, and extragalactic point sources were masked. Masked pixels and missing pixels due to scan strategies were inpainted and finally the coarse and fine angular scales were combined in spherical harmonic space.
Within the frequency range of 45 to 2300 MHz free-free emission is not completely negligible across the full sky and so it must be accounted for in our analysis. We tested two approaches for handling the free-free emission. The first approach removes a free-free template at each frequency and performs a least squares regression between the synchrotron model and the empirical data at each pixel. The second approach separates the synchrotron and free-free components with a Bayesian parametric algorithm fgbuster, commonly adopted by the community for CMB foreground cleaning.
The spectral index and curvature maps were used to predict diffuse Galactic synchrotron emission at 159, 820, 2303 and 2326 and 11000 MHz using the Haslam 408 MHz all-sky map as a synchrotron amplitude template. We ad hoc chose these five frequencies as they span a wide range of interest for both the 21 cm and CMB communities. Moreover, these data sets have not been included in our fit, due to their partial sky coverage. We therefore leverage this to further test and validate the models obtained with the two approaches.
Through this empirical evaluation, it is evident that the combination of free–free template subtraction and a least-squares fitting procedure yielded better results than the alternative approach based on component separation, the latter being more strongly affected by instrumental systematics arising from calibration errors, which in turn led to an underestimation of the associated uncertainties and larger $\chi^{2}$.
We plan to devote future work to overcoming this limitation, by fitting spectral parameters together with calibration errors and offsets within the same framework, employing novel methodologies recently described in the literature (bayes; beyondplanck). A joint fitting approach will enable a proper assessment of the anti-correlation we see between the least-squares parametric spectral index and curvature values, as this may be physical or may simply be due to degeneracies within our fit.
We compare our emission models with those of the pysm3 and GSM. At the test frequencies, we evaluate that the least squares parametric fit provides the closest model to the data at 159, 820 and 2303 MHz, coming second to the GSM at 2326 MHz. Our least-squares parametric model for synchrotron emission provides the most reliable sky model between 45 and 2300 MHz as it consistently predicts the sky temperature to accuracies of around 20 percent on average whilst over methods jump between 10 and 70 per cent in terms of average accuracies. We, therefore, recommend the community to use the templates obtained with the parametric model as it shows higher level of accuracy. By 11 GHz the model starts to show significant deviations from the empirical QUIJOTE data and so this model is not yet complete. However, the power of the parametric framework presented in this paper is that as new data within the 50 to 10000 MHz (synchrotron and free-free dominated) frequency regime become publicly available (e.g. C-BASS (mike)) this sky model can and will be updated and improved.
Acknowledgments
We thank: Brandon Hensley, Susan Clark, Paddy Leahy, Mathieu Remazeilles, Shamik Ghosh, Jacques Delabrouille, Roke Cepeda-Arroita and the Pan-Experiment Galactic Science group for the comments and fruitful discussion and suggestions on synchrotron modelling. We also thank the excellent resource of the LAMBDA Legacy Archive  [^3] https://lambda.gsfc.nasa.gov/ which houses the majority of the sky maps used in this work.
GP acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title “SHIFT” – CUP 55723062008 - Grant Assignment Decree No. 962 adopted on June 30th 2023 by the Italian Ministry of Ministry of University and Research (MUR).
Some of the results in this paper have been derived using the following healpy and HEALPix packages.
Data Availability
We publicly release the code and the scripts used produce the results presented in this paper at https://github.com/giuspugl/fitting_synchrotron. At the same repository, it is possible to download the maps of $\beta_{s}$ and $c_{s}$ described in this work.
Appendix A Empirical data
In Fig. 9, 11, 10, 12, we plot respectively the data sets employed in respectively NC, NF, SC and SF cases used to estimate the synchrotron spectral index and the curvature parameter.
We report in Table 3 the offset values that have been removed to each maps employed before the parameter fitting. We show in Figure 13 the distribution of the estimated offsets together with the maps of the offsets (Figure 14) estimated in a lower pixel grid $N_{\rm side}=8$ .
Figure 9: Frequency maps at 5 resolution employed in the fit of $\beta_{s}$ and $c_{s}$ across the Northern hemisphere. Color scale is log10(K).
Figure 10: Frequency maps at 5 resolution employed in the fit of $\beta_{s}$ and $c_{s}$ across the Southern hemisphere. Color scale is log10(K).
Figure 11: The OVRO-LWA maps at 1 resolution employed in the fit of $\beta_{s}$ and $c_{s}$ across the Northern hemisphere. Color scale is log10(K).
Figure 12: Frequency maps at 1 resolution employed in the fit of $\beta_{s}$ and $c_{s}$ across the Southern hemisphere. Color scale is log10(K).
Figure 13: Histogram distributions of the offset values in Kelvin; the mean (solid black line) and standard error on the mean (dashed black line) are used as the map offset value and uncertainty.
Figure 14: Maps of the offset values calculated within each super pixel for each frequency map. Color scale is in K.
Survey Freq. (MHz) Offset (K)
Maipu/MU 45 $2860\pm 94$
LWA1 50 $2290\pm 70$
LWA1 60 $1608\pm 42$
LWA1 70 $1127\pm 26$
LWA1 74 $1026\pm 22$
LWA1 80 $845\pm 19$
Landecker 150 $84\pm 6$
GMIMS-HBN 1383 $0.39\pm 0.02$
GMIMS-HBN 1418 $0.32\pm 0.01$
GMIMS-HBN 1456 $0.31\pm 0.01$
GMIMS-HBN 1487 $0.31\pm 0.1$
GMIMS-HBN 1499 $0.31\pm 0.1$
GMIMS-HBN 1521 $0.33\pm 0.01$
GMIMS-HBN 1614 $0.27\pm 0.01$
GMIMS-HBN 1625 $0.24\pm 0.01$
GMIMS-HBN 1660 $0.26\pm 0.01$
GMIMS-HBN 1700 $0.22\pm 0.01$
GMIMS-HBN 1712 $0.022\pm 0.01$
GMIMS-STAPS 1324 $0.83\pm 0.01$
GMIMS-STAPS 1349 , $0.77\pm 0.01$
GMIMS-STAPS 1374 $0.73\pm 0.01$
GMIMS-STAPS 1456 $0.63\pm 0.01$
GMIMS-STAPS 1524 $0.57\pm 0.01$
GMIMS-STAPS 1609 $0.49\pm 0.01$
GMIMS-STAPS 1628 $0.49\pm 0.01$
GMIMS-STAPS 1700 $0.45\pm 0.01$
GMIMS-STAPS 1749 $0.43\pm 0.01$
GMIMS-STAPS 1770 $0.41\pm 0.01$
Table 3: The calculated offset values for each of the single-dish empirical data sets used in this work.
Appendix B Least Squares fitting
Figures 15 and 16 show results of the least square parametric fit for the synchrotron spectral index at 45 MHz, curvature and $f$ amplitude parameter for each of the four fits. The northern and southern fits are presented on the same plots.
Figure 15: Results of the northern and southern coarse (5) fit; the first row shows sky maps of the fitted parameters, the second presents the uncertainty maps for each of the parameters and the third row displays the per-pixel reduced $\chi^{2}$.
Figure 16: Results of the northern and southern fine (1) fit; the first row shows sky maps of the fitted parameters, the second presents the uncertainty maps for each of the parameters and the third row displays the per-pixel reduced $\chi^{2}$.
Appendix C fgbuster estimates
In Figure 17, we show the estimated parameters obtained with fgbuster . The NC-SC, (as well as NF-SF) have been combined together into 4 maps of the estimated parameters shown in the top (bottom) panel of Figure 17, with unobserved pixels set to NaN values.
Figure 17: Spectral parameters estimated with fgbuster. (top) and (bottom) panels refer to the parameters estimated respectively for the $5\deg$ and $1\deg$ resolution North and South maps combined into a single map.